if cosx + cosy=4/5 and cosx - cosy=2/7 prove that 14tan(x-y/2)+5cot(x+y/2)=0

Grade:11

3 Answers

Lab Bhattacharjee
8 years ago
\text{Prosthaphaeresis Formulas}\\ \cos x+\cos y=2\cos\dfrac{x+y}2\cos\dfrac{x-y}2\\ \text{ and } \cos x-\cos y=-2\sin\dfrac{x+y}2\sin\dfrac{x-y}2\text{Using Prosthaphaeresis Formulas}\\ \cos x+\cos y=2\cos\dfrac{x+y}2\cos\dfrac{x-y}2\\ \text{ and } \cos x-\cos y=-2\sin\dfrac{x+y}2\sin\dfrac{x-y}2
girish
8 years ago
adding both you get cosx subtracting cosy so we also got sinx and siny if we use half angle formulas we can write sinx in terms of tanx/2 so we got that too so we can solve
Gowri sankar
8 years ago
This question solve using half angle formulas we can write sinx in terms of tanx/2 so we got that then solve this problem.